OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). sin 2x - sin 4x = 0 I NEED HELP I DON'T UNDERSTAND THIS CRAP.

OpenStudy (anonymous):

srry dont really know

OpenStudy (anonymous):

im only in geomerty

OpenStudy (anonymous):

how helpful of you!

OpenStudy (anonymous):

is it multipulchoise

OpenStudy (lukecrayonz):

sin(2x) + sin(4x) = 0 sin(2x) + 2sin(2x)cos(2x) = 0 sin(2x)[1 + 2cos(2x)] = 0 sin(2x) = 0 or 2cos(2x) = -1 -> cos(2x) = -1/2 when sin(2x) = 0 2x = 0, pi, 2pi, 3pi x = 0, pi/2, pi, 3pi/2 when cos(2x) = -1/2 2x = 2pi/3, 4pi/3, 8pi/3, 10pi/3 x = pi/3, 2pi/3, 4pi/3, 5pi/3 so x = 0, pi/2, pi, 3pi/2 , pi/3, 2pi/3, 4pi/3, 5pi/3 @jim_thompson5910 is thsi correct?

OpenStudy (anonymous):

sin(2x) = sin(4x) sin(2x) = sin(2*2x) sin(2x) = 2sin(2x)cos(2x) 2sin(2x)cos(2x) - sin(2x) = 0 sin(2x) * (2cos(2x) - 1) = 0 sin(2x) = 0 or cos(2x) = 1/2 x = 0 + n•π/2, π/6 + n•π, 5π/6 + n•π x = 0, π/6, π/2, 5π/6, π, 7π/6, 3π/2, 11π/6, 2π

OpenStudy (anonymous):

As the king I demand you take your sluttish Profile Picture off and put something more holy. What would your parents say.

OpenStudy (anonymous):

a. π/6 , π/2 , 5π/6 , 7π/6 , 3π/2, 11π/6 b. 0, π/6 , π/2 , 5π/6 , π, 7π/6 , 3π/2, 11π/6 c. 0, 2π/3, 4π/3 d. 0, π/3, 2π/3, π, 4π/3, 5π/3

OpenStudy (anonymous):

as reaper i sujuest u shh or die

OpenStudy (anonymous):

My lord, You people can't take a joke.

OpenStudy (anonymous):

i was jst kiddn

OpenStudy (anonymous):

@AFleming42 dose that help u at all??

OpenStudy (anonymous):

Oh, I wasn't that is a very sluttish picture. Parents you watch what there kids wear these days.

OpenStudy (anonymous):

that was a graduation party

OpenStudy (anonymous):

How did you get that answer @Reaper534 ?

OpenStudy (anonymous):

by the power of google hahaha

OpenStudy (anonymous):

Alright well I'll take google's word on it and hope for the best!

OpenStudy (anonymous):

thanks mucho(:

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

Glad i could help.

jimthompson5910 (jim_thompson5910):

Lukecrayonz you are correct, so that would make the answer choice B

OpenStudy (anonymous):

In that case thank you again @jim_thompson5910 !

jimthompson5910 (jim_thompson5910):

didn't do much, Lukecrayonz did all the work really

OpenStudy (anonymous):

good job @Lukecrayonz (: <3

OpenStudy (lukecrayonz):

thanks cutie can i get your number

OpenStudy (anonymous):

message me(;

OpenStudy (anonymous):

ernies back